# Machine Learning for Event Reconstruction in the CMS Phase-2 High Granularity Calorimeter Endcap

### v1

@@ 1 Introduction
Engineering view of one HGCAL endcap. Source: [4]
@@
In total, the calorimeter will feature about 6 million sensors, each capable of measuring both energy deposition and timing. This combination of fine granularity and precise timing represents a novel approach to calorimetry, enabling improved pileup rejection and particle identification, while preserving excellent energy resolution under HL-LHC conditions. Fully exploiting these capabilities requires the development of advanced reconstruction algorithms tailored to this new detector concept.
@@ 4 Electron reconstruction
Electrons produced in proton-proton collisions radiate bremsstrahlung photons while traversing the CMS tracker, leading to spread-out energy deposits in the calorimeter, an example of which is shown in Figure 4. To recover these photons and maintain good electron momentum resolution, CMS employs superclustering [9], which combines calorimeter clusters associated with the parent electron and seeds track reconstruction with the Gaussian Sum Filter algorithm [10].
@@ 5 Hadron energy regression
For hadronic tracksters, the energy resolution of the calorimeter can be considerably improved with machine-learning based regression. Using a Graph Neural Network, fed with all the reconstructed hits of a test beam setup with charged pions, it was shown that a significant energy resolution improvement on hadronic showers was possible [13], as the neural network can learn the structure of the shower and partly compensate for energy leakage.

### v2

@@ 1 Introduction
Simulated event display at 200 pileup, showing particle trajectories in the tracker (green) and energy deposits in the calorimeter endcaps (blue). Source: CMS Collaboration
@@
In total, the calorimeter will feature about 6 million sensors, each capable of measuring both energy deposition and timing, with occupancy ranging from 60% in the front of the EM section down to less than 1% in the rear scintillator sensors [5]. This combination of fine granularity and precise timing represents a novel approach to calorimetry, enabling improved pileup rejection and particle identification, while preserving excellent energy resolution under HL-LHC conditions. Fully exploiting these capabilities requires the development of advanced reconstruction algorithms tailored to this new detector concept.
@@ 4 Electron reconstruction
Electrons produced in proton-proton collisions radiate bremsstrahlung photons while traversing the CMS tracker, leading to spread-out energy deposits in the calorimeter. To recover these photons and maintain good electron momentum resolution, CMS employs superclustering [9], which combines calorimeter clusters associated with the parent electron and seeds track reconstruction with the Gaussian Sum Filter algorithm [10].
@@ 5 Hadron energy regression
For hadronic tracksters, the energy resolution of the calorimeter can be considerably improved with machine-learning based regression. Using a Graph Neural Network, fed with all the reconstructed hits of a test beam setup with charged pions (without pileup), it was shown that an improvement in the energy resolution of hadronic showers up to a factor 2 was possible [12], as the neural network can learn the structure of the shower and partly compensate for energy leakage.